OpenStudy (anonymous):

Leroy opened a savings account 18 years ago with a deposit of $2,651.43. The account has an interest rate of 2.8% compounded twice a year. How much interest has Leroy earned?

OpenStudy (anonymous):

\[2,651.34\times (1+\frac{0.028}{2})^{36}\] is his current balance

OpenStudy (anonymous):

wow your perfect please move to my house be my mentor!!lmao

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=2%2C651.34 \times+%281%2B\frac{0.028}{2}%29^{36}

OpenStudy (anonymous):

that didn't work, copy and paste looks like current balance is about $4373.53 subtract to get the total interest

OpenStudy (anonymous):

thank, but i don't think i can move today

OpenStudy (anonymous):

You can teleport no biggie

OpenStudy (anonymous):

Was the answer to this problem @satellite73 $1,722.25